Circular No. 9102

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 232P/2009 W1 = 1999 XO_{188} (HILL) As first announced on CBET 2083, F. Manca (Bosisio Parini, Italy) has

suggested that comet P/2009 W1 (cf. IAUCs 9095, 9097) was observed at its previous perihelion passage as 1999 XO_{188}, the observations of which (all by LINEAR, from 1999 Dec. 12 and 15 and 2000 Jan. 2, with the comet’s magnitude given then as ∼ 18–19) are given on MPS 9249 and 82876. An orbit was published on MPO 50446. Additional Mt. Lemmon observations from 2008 Nov. 8 and an orbital linkage are given on MPEC 2009-Y21. This comet has been assigned the permanent number designation 232P.

COMET C/2009 Y1 (CATALINA) As first announced on CBET 2084, an apparently asteroidal object

reported by R. E. Hill (discovery observation tabulated below) using the Catalina Sky Survey 0.68-m Schmidt telescope, after being added to the Minor Planet Center’s ‘NEO Confirmation Page’, was found to be cometary by W. H. Ryan and E. V. Ryan at the Magdalena Ridge Observatory, when observations on Dec. 18.4 and 20.5 UT with the 2.4-m f/8.9 reflector showed a faint tail in p.a. 260^{o}. Hints of cometary activity on the western side of the nuclear condensation were also suspected by H. Sato (remotely using an RAS Observatory 0.25-m f/3.4 reflector near Mayhill, NM, U.S.A.) on Dec. 18.4 and by L. Buzzi (Schiaparelli Observatory, 0.60-m f/4.6 reflector) on Dec. 20.2.

2009 UT α_{2000} δ_{2000} Mag. Observer
Dec. 17.51495 12^{h}10^{m}29.^{s}85 +44^{o}28′31′′.3 19.4 Hill

Further astrometry and an ephemeris were given on MPEC 2009-Y33. The following revised parabolic orbital elements are taken from MPC 67973.

T = 2011 Jan. 27.935 TT  ω = 126.^{o}887
Ω = 160.306 }2000.0
q = 2.53705 AU  i = 107.326

2009 YS_{6}

An asteroidal object discovered via the 1.5-m reflector of the Mt. Lemmon survey on Dec. 17 (red mag ∼ 20) has a retrograde orbit (cf. MPEC 2009-Y35; T ∼ Feb. 2011, q ∼ 1.3 AU, P ∼ 12 yr, i ∼ 144^{o}). Additional astrometry is desired.

2009 December 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green